OpenStudy (anonymous):

If f(x) = −5x + 1, find f(−3).

hero (hero):

Hint x = -3

OpenStudy (anonymous):

We can plug -3 where it says x: f(-3)=-5*(-3)+1 How do we continue now?

OpenStudy (anonymous):

-5(3)+1=-15+1

OpenStudy (anonymous):

@nia_maepenix not -15...should be -5(-3) + 1

OpenStudy (anonymous):

@nia_maepenix we are looking for f(-3), not f(3)

OpenStudy (anonymous):

sorry -5(-3)

OpenStudy (anonymous):

15

OpenStudy (anonymous):

@ivandelgado Are you there? Is this useful? f(-3)=-5*(-3)+1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its 15

OpenStudy (anonymous):

Be careful, -5*(-3)=15, but we have to add 1, right? Look at the original equation.

OpenStudy (anonymous):

its not 15 ?

OpenStudy (anonymous):

No, f(-3)=-5*(-3)+1=15+1=16 So f(-3)=16 Do you agree with that ?

OpenStudy (anonymous):

Thanks for that medal! :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Any doubt about this problem? Feel free to ask!